# Rotational mass of anisotropic neutron stars within Rastall gravity

### v3

@@ 1 Introduction
Berti et al. [14] conducted a comparative analysis of the HT approximation, the Cook–Shapiro–Teukolsky (CST) solutions [17], and the Manko–Miel-ke–Sanabria-Gómez (MMSG) formalism [18]. Note that the CST framework accommodates both slowly and rapidly rotating compact stars. In the comparison between the HT approximation and the CST models, configurations were selected such that the gravitational mass and angular momentum were matched. Their results demonstrate that the HT formalism provides remarkably accurate predictions for the corotating and counterrotating radii, $R_{\pm}$, of the innermost stable circular orbit, with deviations less than 1 % even for the most rapidly rotating millisecond pulsars. When the HT approximation is compared to the MMSG formalism—specifically in the exterior vacuum spacetime—it also demonstrates better overall performance and higher accuracy.
@@ 2.2 Nuclear model of the EOS
where ${W_{\mu\nu}}={\partial_{\mu}\omega_{\nu}-\partial_{\nu}\omega_{\mu}}$ and ${\vec{R}_{\mu\nu}}={\partial_{\mu}\vec{\rho}_{\nu}-\partial_{\nu}\vec{\rho}_{\mu}.}$ Here $\sigma$ and $\omega$ are isoscalar mesons, and $\vec{\pi}$ and $\vec{\rho}$ are isovector meson.
@@ 3 Numerical results and discussion
Another important differential equation is Eq. (LABEL:omega), which contributes in giving the value of factor $\bar{\omega}$ and $\frac{d\bar{\omega}}{dr}$ in other differential equations. Since the value of $\bar{\omega}_{c}$ as the initial value of $\bar{\omega}$ can be chosen arbitrarily, we choose 240 $s^{-1}$ in this work. It has to be highlighted that this value is much less than the value used in Ref. [4], i.e. $\bar{\omega}$ = 3000 s-1. Choosing our value of $\bar{\omega}_{c}$ is also relatively wise since it is more suitable for HT formalism. We have to consider that $\bar{\omega}(r)$ increases as $r$ increases. Although we have a fact that according to Ref. [14], the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars, putting $\bar{\omega}_{c}$ = 3000 s-1 in the calculation is potential to be awful, for the HT formalism is basically constructed for slowly rotating relativistic stars. So our value of $\bar{\omega}_{c}$ is safer than the one in Ref. [4]. Comparing to Ref. [7] and Ref. [16] which chose $\bar{\omega}_{c}=0$, our choice of $\bar{\omega}_{c}$ is also more realistic for general rotating NS, for their choice tends to only agree the NSs with very slow rotation.

### v5

@@ 1 Introduction
Berti et al. [14] conducted a comparative analysis of the HT approximation, the Cook–Shapiro–Teukolsky (CST) solutions [17], and the Manko–Mi-elke–Sanabria-Gómez (MMSG) formalism [18]. Note that the CST framework accommodates both slowly and rapidly rotating compact stars. In the comparison between the HT approximation and the CST models, configurations were selected such that the gravitational mass and angular momentum were matched. Their results demonstrate that the HT formalism provides remarkably accurate predictions for the corotating and counterrotating radii, $R_{\pm}$, of the innermost stable circular orbit, with deviations less than 1 % even for the most rapidly rotating millisecond pulsars. When the HT approximation is compared to the MMSG formalism—specifically in the exterior vacuum spacetime—it also demonstrates better overall performance and higher accuracy.
@@ 2.2 Nuclear model of the EOS
${\text{where}}\>{W_{\mu\nu}}={\partial_{\mu}\omega_{\nu}-\partial_{\nu}\omega_{\mu}}$ and ${\vec{R}_{\mu\nu}}={\partial_{\mu}\vec{\rho}_{\nu}-\partial_{\nu}\vec{\rho}_{\mu}.}$ Here $\sigma$ and $\omega$ are isoscalar mesons, and $\vec{\pi}$ and $\vec{\rho}$ are isovector meson.
@@ 3 Numerical results and discussion
The procedure to obtain $\bar{\omega}(r)$ and $\frac{d\bar{\omega}}{dr}$ is presented in Ref. [16]. Note that Rastall parameter $\lambda$ does not affect the calculations of $\bar{\omega}(r)$ and $\frac{d\bar{\omega}}{dr}$. Since the value of $\bar{\omega}_{c}$ as the initial value of $\bar{\omega}$ can be chosen arbitrarily, we choose 240 $s^{-1}$ in this work. It has to be highlighted that this value is much less than the value used in Ref. [4], i.e. $\bar{\omega}_{c}$ = 3000 s-1. Choosing our value of $\bar{\omega}_{c}$ is also relatively wise since it is more suitable for HT formalism. We have to consider that $\bar{\omega}(r)$ increases as $r$ increases. Although we have a fact that according to Ref. [14], the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars, putting $\bar{\omega}_{c}$ = 3000 s-1 in the calculation is potential to be awful, for the HT formalism is basically constructed for slowly rotating relativistic stars. So our value of $\bar{\omega}_{c}$ is safer than the one in Ref. [4]. Comparing to Ref. [7] and Ref. [16] which chose $\bar{\omega}_{c}=0$, our choice of $\bar{\omega}_{c}$ is also more realistic for general rotating NS, for their choice tends to only agree the NSs with very slow rotation.
@@ Acknowledgements
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, M. Lawrence Pattersons used ChatGPT (OpenAI) in order to improve the language and readability of the manuscript. After using this tool, he reviewed and edited the content as needed and take full responsibility for the content of the publication.